An aeroplane is flying horizontally with a velocity of 600 km/h and at a height of 1960 m. When it is vertically at a point A on the ground a bomb is released from it. The bomb strikes the ground at point B. The distance AB is

Grade:

2 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Convert velocity in horizontal direction, Ux into m/s:
Ux = 600x1000/3600 = 500/3 m/s
Velocity of the bomb during the flight is constant due to inertia. Therefore,
Initial velocity in vertical direction, Uy = 0
Height = 1960 m
Step 2: Find time ‘t’ using the equation [s = ut + 1/2at^2]
Replace distance ‘s’ by height ‘h’ and acceleration ‘a’ by gravity ‘g’
h = 1/2gt^2
t = √(2h)/g = √(2x 1960/9.8) = 20 sec.
Step 3: Find the distance traversed by the bomb in horizontal direction:
Distance AB = vx t = 500/3 x 20 = 10/3 x 10^3m = 3.333 km.
Thanks and Regards
Shaik Aasif
askIItians faculty
Kushagra Madhukar
3 years ago
Dear student,
Please find the answer to your problem below.
 
Step 1: Convert velocity in horizontal direction, Ux into m/s:
Ux = 600x1000/3600 = 500/3 m/s
Velocity of the bomb during the flight is constant due to inertia. Therefore,
Initial velocity in vertical direction, Uy = 0
Height = 1960 m
Step 2: Find time ‘t’ using the equation [s = ut + 1/2at^2]
Replace distance ‘s’ by height ‘h’ and acceleration ‘a’ by gravity ‘g’
h = 1/2gt^2
t = √(2h)/g = √(2x 1960/9.8) = 20 sec.
Step 3: Find the distance traversed by the bomb in horizontal direction:
Distance AB = vx t = 500/3 x 20 = 10/3 x 10^3m = 3.333 km.
 
Hope it helps.
Thanks and regards,
Kushagra